OpenStudy (anonymous):

What is the difference between the sum of all even numbers and the sum of all odd numbers from 0 through 1000?

OpenStudy (anonymous):

even \[n^2+n\] odd \[n^2\]

OpenStudy (aravindg):

sum of even numbers till 1000 - sum of odd numbers till 1000 = (2 + 4 + 6 + ... + 1000) - (1 + 3 + 5 + ... + 999) =2 + 4 + 6 + ... + 1000 - 1 - 3 - 5 - ... - 999 =2 - 1 + 4 - 3 + 6 - 5 + ... + 1000 - 999 =(2 - 1) + (4 - 3) + (6 - 5) + ... (1000 - 999) Does this help ??

OpenStudy (anonymous):

Not really....

hero (hero):

^Yes, @AravindG is on the right track

OpenStudy (aravindg):

why not ? what doubt do you have in my working?

hero (hero):

The difference increases by 5 every 10 units

OpenStudy (anonymous):

So it's 500? That's hard to understand...

OpenStudy (anonymous):

\[1000=2k-1 ,1000=2k \implies k=500 \] difference \[n^2+n-n^2=n\] 500

OpenStudy (anonymous):

Ah. Thanks!

hero (hero):

Great job everyone